# SpikySpace: A Spiking State Space Model for Energy-Efficient Time Series Forecasting

### v1

@@ Abstract
Time-series forecasting often operates under tight power and latency budgets in fields like traffic management, industrial condition monitoring, and on-device sensing. These applications frequently require near real-time responses and low energy consumption on edge devices. Spiking neural networks (SNNs) offer event-driven computation and ultra-low power by exploiting temporal sparsity and multiplication-free computation. Yet existing SNN-based time-series forecasters often inherit complex transformer blocks, thereby losing much of the efficiency benefit. To solve the problem, we propose SpikySpace, a spiking state-space model (SSM) that reduces the quadratic cost in the attention block to linear time via selective scanning. Further, we replace dense SSM updates with sparse spike trains and execute selective scans only on spike events, thereby avoiding dense multiplications while preserving the SSM’s structured memory. Because complex operations such as exponentials and divisions are costly on neuromorphic chips, we introduce simplified approximations of SiLU and Softplus to enable a neuromorphic-friendly model architecture. In matched settings, SpikeSpace reduces estimated energy consumption by 98.73% and 96.24% compared to two state-of-the-art transformer based approaches, namely iTransformer and iSpikformer, respecitvely. In standard time series forecasting datasets, SpikeSpace delivers competitive accuracy while substantially reducing energy cost and memory traffic. As the first full spiking state-space model, SpikySpace bridges neuromorphic efficiency with modern sequence modeling, marking a practical and scalable path toward efficient time series forecasting systems.
I Introduction
Edge-deployed time-series analytics systems, such as urban traffic management, industrial condition monitoring, and on-device wearable sensing, are driving a need for time-series forecasting methods that operate on the edge. Due to privacy, reliability or bandwidth constraints, these tasks must produce real-time predictions locally without relying on the cloud. While deep learning methods including RNNs, LSTMs, and GRUs or more recently transformer-based structures can achieve high accuracy, they still remain much energy hungry according to complex multiplication and exponentiation operations. To meet the energy and latency demands, Spiking Neural Networks (SNNs), which adopts the event-driven paradigm and only adapt simplier addition and shifting operations, has emerged as a promising direction towards edge intelligence [1]. This mechanism allows compute and memory activity to scale with signal changes rather than sequence length, enabling low-latency operation on edge devices. Furthermore, SNNs possess an intrinsic ability to model temporal dynamics through membrane potentials and spike timing, making it well-suited for streaming time-series forecasting in scenarios like sensor-based monitoring and edge-level control, where both energy efficiency and real-time responsiveness are critical.
However, most recent SNN works focus on better performance on time series tasks instead of energy efficiency. Works like SpikeTCN [2], iSpikformer [2], and TS-LIF [3] have achieved desirable performance, while they still rely on high-cost encoders or the transformer architecture, whose attention mechanism remains energy-intensive due to frequent memory access and global communication, even under spiking implementations. Although these previous works achieved desirable performance, their improvements often come at the cost of increased energy consumption. While pursuing higher accuracy is important, edge-deployed forecasting systems must ultimately balance accuracy against deployability. These limitations highlight the need for a more energy-efficient neuromorphic time-series model that achieves competitive predictive performance under stringent energy and memory constraints.
In this work, we propose SpikySpace, a novel spiking model for time series forecasting that addresses the imbalance between accuracy and energy efficiency in prior spiking approaches. SpikySpace adopts an event-driven state space architecture that combines selective spike-based updates with continuous time dynamics, allowing it to model temporal dependencies efficiently under strict power and memory constraints. Through the joint optimization of structural sparsity and spike-driven state transitions, SpikySpace achieves accuracy comparable to or better than existing models while greatly reducing power consumption.
Specifically, in SpikySpace, we leverage the state-space model (SSM) [4, 5] formulation to replace the high computational cost of the transformer while maintaining its strong modeling capacity. Instead of relying on sequential hidden-state updates as in RNNs, LSTMs, and GRUs, or on dense self-attention and large context buffers as in transformers, SSMs represent temporal dynamics through compact latent states updated via linear-time recurrences. This formulation allows information to propagate over long horizons with $O(L)$ complexity, significantly reducing both computation and memory costs compared to the $O(L^{2})$ attention in transformers and avoiding the sequential bottleneck of recurrent networks. By embedding this formulation into an event-driven spiking framework, SpikySpace inherits the long-range dependency modeling of SSMs while maintaining the sparse and energy-efficient computation of SNNs. Consequently, SpikySpace matches transformer-level forecasting performance with far lower energy consumption, making it well-suited for deployment on neuromorphic or edge platforms.
Developing SSM-based SNNs remains challenging due to the limited explore of fully spiking state updating as well as the complex activation functions involved. SpikySpace optimize the nonlinearities such as SiLU and Softplus employed in gating and readout in modern SSMs for neuromorphic hardware compatibility. We first develop SNN-compatible approximations for Softplus and SiLU with power-of-two components and linear transformations, namely PTsoftplus and PTSiLU. We theoretically prove our proposed components are continously differentiable and act close to the original functions. This design allows the activation to be implemented through efficient bit-shifting operations, removing costly divisions and exponentials while preserving smooth gating behavior. For training, we first learn a quantized ANN in the continuous domain using the proposed PTsoftplus and PTSiLU approximations under low-bit constraints for activations. We then perform a deterministic ANN to SNN conversion that maps the quantized activations, gates, and state updates to spike events. The resulting SpikySpace is a spiking forecaster suitable for resource-constrained edge deployment. While several prior efforts have attempted to integrate SNNs and SSMs [6, 7, 8, 9], their formulations still partly rely on dense multiplications and retain complicated activations such as Softplus or SiLU, limiting their spiking sparsity and hardware efficiency.
Evaluated on four multivariate forecasting datasets, SpikySpace achieves competitive or superior accuracy while substantially reducing parameter size and energy consumption. Specifically, on datasets like Metr-la and Electricity [10, 11], SpikySpace attains average $R^{2}$ scores of $0.778$ and $0.992$, surpassing the accuracy of the previous best-performing SNN SpikeSTAG by 3.3% and 0.7%, respectively, while consuming only 2.7% of its energy. SpikySpace also achieves the energy savings by $26.58\times$ over iSpikformer and $78.9\times$ over iTransformer, without compromising accuracy stability. We further conduct ablation studies on the proposed PTsoftplus and PTSiLU and spiking timesteps, verifying their contribution and justify our parameter selection. Mechanistically, SpikySpace retains information over long horizons through its explicit latent state, while event-driven spikes naturally skip redundant updates. The result shows that structured spiking state evolution enables efficient and scalable time-series forecasting on neuromorphic and edge devices.

### v2

@@ Abstract
Time-series forecasting in domains like traffic management and industrial monitoring often requires real-time, energy-efficient processing on edge devices with limited resources. Spiking neural networks (SNNs) offer event-driven computation and ultra-low power and have been proposed for use in this space. Unfortunately, existing SNN-based time-series forecasters often use complex transformer blocks. To address this issue, we propose SpikySpace, a spiking state-space model (SSM) that reduces the quadratic cost in the attention block to linear time via spiking selective scanning. Further, we introduce PTsoftplus and PTSiLU, two efficient approximations of SiLU and Softplus that replace costly exponential and division operations with simple bit-shifts. Evaluated on four multivariate time-series benchmarks, SpikySpace outperforms the leading SNN in terms of accuracy by up to 3.0% while reducing energy consumption by over 96.1%. As the first fully spiking state-space model, SpikySpace bridges neuromorphic efficiency with modern sequence modeling, opening a practical path toward efficient time series forecasting systems. Our code is available at https://anonymous.4open.science/r/SpikySpace
Machine Learning, ICML
1 Introduction
Analytics systems deployed at the edge, such as urban traffic management, industrial monitoring, and on-device sensing, demand forecasting methods that function locally (Lv et al., 2022; Reis and Serôdio, 2025; Kashpruk et al., 2023). Driven by strict limits on power, privacy, and bandwidth, these applications require predictions in real time without reliance on the cloud. While accurate, deep learning models rely on dense matrix multiplications that consume substantial energy, making them difficult to deploy on devices with limited power reserves. Spiking neural networks (SNNs) offer a promising alternative (Han et al., 2022). By replacing complex multiplications with discrete additions driven by events, SNNs enable computation to scale sparsely with signal changes (Yamazaki et al., 2022). This mechanism is naturally suited for processing streaming data, where efficiency and responsiveness are essential.
Despite these advantages, recent advancements in SNNs have yet to translate theoretical efficiency into practice. Works like SpikeTCN (Lv et al., 2024), iSpikformer (Lv et al., 2024), and TS-LIF (SHIBO et al., 2025) have achieved desirable performance, while they still rely on high-cost encoders or the transformer architecture. Even with spikes, the underlying mechanism of self-attention exhibits quadratic complexity with respect to the length of the sequence, which causes the costs of memory access and communication to explode over time. This fundamentally undermines the energy benefits of event-driven computation, highlighting the need for a more efficient architectural alternative.
To bridge this gap, we propose SpikySpace, a novel spiking architecture that leverages State Space Models (SSMs) (Gu et al., 2021; Gu and Dao, 2024) to achieve high accuracy under strict energy constraints. SpikySpace adopts an architecture based on event-driven state spaces that combines selective updates driven by spikes with continuous time dynamics. Unlike recent SNNs constrained by the quadratic complexity of attention mechanisms or the sequential bottlenecks of RNNs, SpikySpace employs compact latent states updated with $O(L)$ complexity. This design eliminates the heavy $O(L^{2})$ computation and memory overhead of attention mechanisms while capturing long-range dependencies more effectively than traditional recurrence. Consequently, SpikySpace achieves competitive accuracy while significantly reducing energy consumption, positioning it as an ideal solution for deployment on resource-constrained neuromorphic and edge platforms.
However, implementing SSMs on neuromorphic hardware remains challenging due to the continuous nature of the state selection process and the complexity of activation functions. Existing efforts to integrate SNNs and SSMs (Shen et al., 2025; Bal and Sengupta, 2024; Zhong et al., 2024; Huang et al., 2025) often compromise efficiency by retaining dense floating point multiplications in the recurrent scan or relying on expensive operations like exponentials. SpikySpace overcomes these limitations through a fully discrete hardware-friendly design. First, we introduce a selective scan mechanism driven by spikes. Unlike standard implementations that update states via continuous matrix multiplications, we constrain the parameters for time scales to powers of two and encode all intermediate activations and recurrent states into sparse spikes. This transforms the core state transitions into efficient bit-shift operations and integer accumulations, free of multiplication. Next, we propose approximations we called PTsoftplus and PTSiLU to handle the complex nonlinearities. They utilize linear transformations and power-of-two components to eliminate costly exponentials and divisions. We will give theoretical proofs of their error bounds, guaranteeing that the surrogate gradients remain stable during backpropagation. To enable effective learning, we employ a quantization-aware training strategy in the continuous domain, followed by a deterministic conversion to spike events. To the best of our knowledge, SpikySpace is the first spiking state-space model to realize fully multiplication-free recurrent dynamics.
Our evaluations on four multivariate forecasting benchmarks demonstrate that SpikySpace maintains high predictive performance while drastically lowering resource usage. Specifically, SpikySpace surpasses the accuracy of the previous best-performing SNN SpikeSTAG (Hu et al., 2025) by 3.0% on dataset Metr-la (Li et al., 2018), and even outperform dense iTransformer (Liu et al., ) on dataset Electricity (Lai et al., 2018). SpikySpace also reduces energy consumption by factors of $26.6\times$ compared to iSpikformer and $78.9\times$ compared to iTransformer, without compromising stability.